OpenStudy (rizags):

Solve for all x such that (2x+1)^2 + 13(2x+1) +27 =0

OpenStudy (jdoe0001):

I'm thinking you may wan to use the quadratic formula

OpenStudy (rizags):

i have to use completing the square

jigglypuff314 (jigglypuff314):

I don't think it is possible to find x by completing the square for this equation. in any case, I was only able to factor it out to 4x^2+30x+41=0 and if you want an answer for x, I think it would come out to (1/4)((root(61))-15) which approximates out to about -5.7026 and/or -1.7974 I hope this helps a little...

OpenStudy (solomonzelman):

(2x+1)^2 + 13(2x+1) +27 =0 factor out 4x^2+4x+1+26x+13+27=0 simplify 4x^2+30x+41=0 you can start completing the square now (2x+7.5)^2+41=0+56.25 see how I am doing this?

OpenStudy (solomonzelman):

now subtract 41 from each side (2x+7.5)^2+41-41=56.25 -41 (2x+7.5)^2=15.25 square both sides 2x+7.5= √(15.25) can you do it from there?

OpenStudy (rizags):

yes, thank you

OpenStudy (solomonzelman):

mention my name for help!

OpenStudy (rizags):

sure

OpenStudy (solomonzelman):

I mean in the future.... Glad I helped you!